OpenStudy (anonymous):

Find the value of 10 times the square root of 2 over the square root of 5..

OpenStudy (nurali):

did you mean simplify??? 10√2 / √5 = (10√2 * √5) / (√5 * √5) = 10√10 / 5 = 2√10

OpenStudy (anonymous):